# Narrowband, Fast, and Autonomous Drone Radio Mapping for Localization

### v1

@@ III Results and Analysis
Table I presents the results from the AERPAW emulator and testbed environments. The first mission shows how quickly the GP can fit a transmitter that is relatively close (125 m) from the drone’s starting location. However, the drone could not resolve a better 10-minute location estimate due to receiver gain saturation. While a consistently high gain is generally desired for communication, corresponding to an ideal channel, in this case, the signal ’clipping’ makes a spatial ’plateau.’ This spatial feature makes resolving a transmitter location nearly impossible when within this region. Seeing this within the emulator, the radius for the auxiliary circle was set expecting to encounter this signal behavior; however, in this case, the auxiliary path was not wide enough to provide sufficient localization information. Additionally, critical points were missed on the northwest side of the auxiliary path due to excessive noise, correctly excluded by the quality-variance filter.
@@
Figure 4: This figure shows a reconstruction of the final Gaussian process estimate for the first and second run, along with the actual rover location shown as a red dot. The green dot shows the final ten-minute estimate, and the blue dot indicates the three-minute estimate. This GP reconstruction uses the final kernel hyperparameters indicated by the vehicle log. Part of this figure is adapted from Fig. 4c of [7].
@@
The second trial, shown in Figure 4(b), gives the team’s best 10-minute result of 27.8 m. The 3-minute result did not provide quite enough time for the drone to establish a proper gradient and travel to the transmitter 330 m from takeoff. However, the drone was able to establish a proper optimal trajectory toward the transmitter and perform the auxiliary path maneuver while staying within the boundary conditions. The auxillary path was wide enough to provide additional localization to the GP, and the quality-variance filter was working appropriately to filter excessive noise. This result shows the ideal behavior of the drone in this situation. To provide a better three-minute GP estimate with this starting routine, a kernel that supplied more information about the signal from this starting distance (as a prior) would need to be considered.
@@
During the third trial, the drone encountered a confluence of issues, which resulted in degraded localization estimation performance. However, the mission was not a failure; the drone was able to produce locations for both timed estimates from a starting distance of 286 m. Log analysis revealed that the quality-variance filter threshold (set with the first sample) was set to exclude any variance above 35. This was different from the second run, which allowed samples with a less strict value of 67. This disallowed many of the first critical direction-finding waypoints from being included within the GP. The drone then generated the auxiliary path due to missing too many sample measurements. From this waypoint path two samples were accepted. These samples greatly increased the accuracy of the location estimate. The drone reestablished an optimal path but was unable to collect any more samples due to the time limit.
@@ IV Conclusion
The results of the three trials show significant robustness in the GP and BO architecture for drone-based location finding. The quality-variance filter identified and removed outlier noise values from the narrowband receiver. Unfortunately, the drone was unable to find the correct filter parameter value from the start of the third trial. This alone shows a necessity for improved performance of the AERPAW emulator and a need for radio data collection. If the emulation noise and the testbed had a reliable digital-twin agreement, the filter parameter could be safely studied and set within the emulator similar to the auxiliary path area. This would save mission time and prevent mission hazardous calibrations.

### v2

@@ III Results and Analysis
Table I presents the results from the AERPAW emulator and testbed environments. The first testbed mission shows how quickly the GP can fit a transmitter that is relatively close (125 m) from the drone’s starting location. However, the drone could not resolve a better 10-minute location estimate due to receiver gain saturation. While a consistently high gain is generally desired for communication, corresponding to an ideal channel, in this case, the signal ’clipping’ makes a spatial ’plateau.’ This spatial feature makes resolving a transmitter location nearly impossible when within this region. Seeing this within the emulator, the radius for the auxiliary circle was set expecting to encounter this signal behavior; however, in this case, the auxiliary path was not wide enough to provide sufficient localization information. Additionally, critical points were missed on the northwest side of the auxiliary path due to noise and correctly excluded by the quality-variance filter.
@@
Figure 4: This figure shows the first and second run reconstructions of the final GP estimates over the search spaces shown in figure 1. The actual rover location is shown as a red dot. The green dot shows the final ten-minute estimate, and the blue dot indicates the three-minute estimate. This GP reconstruction uses the final kernel hyperparameters indicated by the vehicle log. Part of this figure is adapted from Fig. 4c of [7].
@@
The second trial, shown in Figure 4(b), gives the best 10-minute result of 27.8 m. The 3-minute result did not provide quite enough time for the drone to establish a proper gradient and travel to the transmitter 330 m from takeoff. However, the drone was able to establish a proper optimal trajectory toward the transmitter and perform the auxiliary path maneuver while staying within the boundary conditions. The auxiliary path was wide enough to provide additional localization to the GP, and the quality-variance filter was working appropriately to filter excessive noise. This result shows the ideal behavior of the drone in this situation. To provide a better three-minute GP estimate with this starting routine, a kernel that supplied more information about the signal from this starting distance (as a prior) would need to be considered.
@@
During the third trial, the drone encountered a confluence of issues, which resulted in degraded estimation performance. However, the mission was not a failure; the drone was able to produce locations for both timed estimates from a starting distance of 286 m. Log analysis revealed that the quality-variance filter threshold (set with the first sample) was set to exclude any variance above 35. This was different from the second run, which allowed samples with a less strict value of 67. This disallowed many of the first critical direction-finding waypoints from being included within the GP. The drone then generated the auxiliary path due to missing too many sample measurements. From this waypoint path, two samples were accepted. These samples greatly increased the accuracy of the location estimate. The drone reestablished an optimal path but was unable to collect any more samples due to the time limit.
@@ IV Conclusion
The results of the three trials show significant robustness in the GP-based BO approach for drone-based localization. The quality-variance filter identified and removed outlier noise values from the narrowband receiver. Unfortunately, the drone was unable to find the correct filter parameter value from the start of the third trial. This alone shows a necessity for bridging the gap between the AERPAW emulator and real-world radio data. If the emulation noise and the testbed had a reliable digital-twin agreement, the filter parameter could be safely studied and set within the emulator, similar to how the auxiliary path area was tested and adopted. This would save mission time and prevent mission-hazardous calibrations. Additionally, by more closely integrating the digital twin emulation/testbed environment, various “black box” pathloss regression estimators for mapping and localization could be trained and trialed within an emulated environment with accurate synthetic radio data.